A monochromatic light illuminates the surface of sodium and the maximum kinetic energy of a photoelectron is 1.2eV.The cutoff wavelength of sodium is 540nm, the wavelength of the incident light is.........


1
Expert's answer
2021-06-18T11:45:56-0400

Given,

The kinetic energy of the photoelectric = 1.2eV

Cut-off wavelength of sodium "(\lambda_o)=540nm = 540\times 10^{-9}m"

Let the wavelength of the incident light be "\lambda."

Now, applying the photoelectric formula,

"eV = \frac{hc}{\lambda}-\frac{hc}{\lambda_o}"

Now, substituting the values,

"\Rightarrow 1.2eV = \frac{1240 eV -nm}{\lambda}-\frac{1240eV-nm}{540nm}"

Now, substituting the values,

"\Rightarrow \frac{1.2eV}{1240 eV-nm}=\frac{1}{\lambda}-\frac{1}{540nm}"


"\Rightarrow \frac{1}{\lambda}=\frac{1.2}{1240nm}+\frac{1}{540nm}"


"\Rightarrow \frac{1}{\lambda}=9.68\times 10^{-4}+18.51\times10^{-4}"


"\Rightarrow \frac{1}{\lambda}=28.19\times 10^{-4}"


"\Rightarrow \lambda = 354.73nm"